The equation of a particle executes s.h.m y=8sin (pi t+pi÷3)find a the amplitude








The period








Maximum velocity

1
Expert's answer
2022-02-06T14:33:53-0500

Given:

"y=8\sin (\pi t+\pi/3)\:\rm cm"


The general equation of SHM is given by

"y=A\sin\left (\frac{2\pi}{T} t+\phi_0\right)"


(a) the amplitude

"A=8\:\rm cm"

(b) the period of oscillation

"T=2\:\rm s"

(c) maximum velocity

"v_{\max}=A\omega=A*2\pi /T\\
=8*2\pi/2=8\pi \:\rm cm/s=25\: cm /s"